OpenStudy (anonymous):

which of the following solutions will have the highest electrical conductivity? A.0.065M Na3PO4 B.0.075M (NH4)2CO3 C.0.10M LiBr D.0.10M NaI E.0.10M KF Can someone please explain it to me.

OpenStudy (aaronq):

in a solution, electrical conductivity is created by ions. The more ions the more electoral conductivity. find which solution produces the most ions

OpenStudy (anonymous):

How would i find out which one processes more ions?

OpenStudy (aaronq):

use the molarity and it's dissociation equation for example: \(NaI\rightarrow Na^+ +I^-\) produces 2 ions 0.10M *2 ions = 0.2 M find moles, the find # of ions Assuming were using the same volume of all solutions: 1L 0.2 M*1L = 0.2 moles -> 0.2 moles * 6.022*10^23 = 1.2044*10^23 ions

OpenStudy (anonymous):

okay thanks

OpenStudy (aaronq):

no problem